OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (-2, -3) with a slope of -6. y = -6x - 15 y = -6x - 20 y = -6x + 15 y = -6x + 20

OpenStudy (anonymous):

I got A

OpenStudy (the_fizicx99):

y-y1=m(x-x1) You'll get your y=mx+b after you plug in your numbers and solve.

OpenStudy (the_fizicx99):

Y - (-3)=-6(x-(-2) Can you solve?

OpenStudy (anonymous):

yeah you do the distributive property and get y + 3 = -6x - 12 And then you subtract 3 from both sides y = -6x - 15 and thats how i got A

OpenStudy (the_fizicx99):

Correct! :p good job

OpenStudy (anonymous):

thx for your help

OpenStudy (the_fizicx99):

My pleasure :}